OpenStudy (anonymous):

for f(x)=3x-7 find: f(2)

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

its beacuse u have to plug in 2 for all of x values in the fucnction thats why it says f(2). f(x)=3x-7 f(2)=3(2)-7 ... replace every 'x' with '2' f(2) = -1

OpenStudy (anonymous):

f(2)= 3(2)-7 F(2)= -1

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

ur welcome :D

OpenStudy (anonymous):

It's confusing that f(2) doesn't mean the same thing as 3(2) here!

OpenStudy (anonymous):

Find the equation of the line you think best represents the average time to brake at the given speed according to the graph: S=6x+60 S=6x-60 S=-6x+60 S=6x-60

OpenStudy (anonymous):

I have the graph, but can not show on here.

OpenStudy (anonymous):

Please give me hint. Thank you.